OpenStudy (anonymous):

Suppose that $70,000 is invented at 9% interest. Find the amount of money in the account after 4 years if the interest is compounded annually. If interest is compounded annually, what is the amount of money after t=4years? $_____

OpenStudy (anonymous):

FV = PV(1+i)^n FV = 70,000(1+0.09)^4 FV = 98810.71

OpenStudy (anonymous):

You are officially my favorite person on here. Think you could help me with 3 more?

OpenStudy (anonymous):

yea sure post em i[

OpenStudy (anonymous):

up*

OpenStudy (anonymous):

ok one sec